Beale Manufacturing Company has a beta of 1, and Foley Industries has a beta of 0.7. The required return on an index fund that holds the entire stock market is 11.5%. The risk-free rate of interest is 7%. By how much does Beale's required return exceed Foley's required return? Round your answer to two decimal places.
Answer #1

required return= risk-free rate +Beta*(market rate- risk-free rate )

required return for Beale=7+1*(11.5-7)=11.5%

required return for Foley=7+0.7*(11.5-7)=10.15%

Hence excess required return=(11.5-10.15)

=1.35%